Use the chain rule to find the derivative of 4V50% + 6x8 Type your answer without fractional or negative exponents. Use sqrt(If f(x) = 7 – x2 5 + x2 find: f() =Evaluate the limit: 9x – 9 lim 2+1 x2 8x + 7 =
Answer #1

Find the derivative of 4 [5x5+6x8 chain rule - Suppose y = f(l) and u= g(x) then df. du du dx y = df ur 4 Suppose y = 45u anf(x) = x ( Clustend . ② af f(x) = 7-X² 5+82 7-X² 5+x2 Quotient rule- a) ule-u. U 2 Suppose u=7-12 and l=5+x2 W = 2x > f(x) =